OpenStudy (zale101):

Integrate:\(\int\limits_{-1}^{1}(1+\sqrt{1-x^2})dx\)

OpenStudy (zale101):

I tried using trig substitution but got lost halfway..

OpenStudy (anonymous):

Actually if you graph this function, you'll know the answer without even finding an anti-derivative.

OpenStudy (zale101):

But, the question doesn't want me to graph it :\

OpenStudy (anonymous):

But anyway, as for \[ \sqrt{1-x^2}~dx \]You would set \(x=\sin t\) \[ \sqrt{1-\sin^2 t} ~d(\sin t) = \sqrt{\cos^2 t} (\cos t)~dt=|\cos t| \cos t~dt \]

OpenStudy (zale101):

\(x+\int\limits_{-1}^{1}(\sqrt{1-sin^2t})*cost dx\) \(x+\int\limits_{x=-1}^{x=1}(cost)*cost dt\) \(x+\int\limits_{x=-1}^{x=1}(cos^2t dt\) cos^2t=(1+cos2t)/2 \(x+\int\limits_{x=-1}^{x=1}(\frac{1+cos^2t}{2} *dt\)

OpenStudy (zale101):

This is what i did

OpenStudy (anonymous):

The limits \(x\in (-1,1)\) translate to \(t\in (-\pi/2,\pi,2\), and so \(|cos t| = \cos t\).

OpenStudy (zale101):

Right, right

OpenStudy (anonymous):

Why do you have the square term?

OpenStudy (zale101):

correction: \(x+\int\limits_{-\pi/2}^{\pi/2}(\frac{1+cos(2t)}{2}\) sorry about that

OpenStudy (anonymous):

Okay, I'm still confused though. Why do you have an \(x\)? The variable of integration goes away when you integrate.

OpenStudy (zale101):

i was thinking of antiderivative of 1 be x?

OpenStudy (anonymous):

\[ \int_{-1}^1dx = 2 \]

OpenStudy (anonymous):

Anyway, I think you can do the rest

OpenStudy (zale101):

I should have applied FTOC before, and consider it more xD

OpenStudy (anonymous):

?

OpenStudy (zale101):

\[2+\frac{1}{2}\int\limits_{-\pi/2}^{\pi/2}1+\cos(2t) dt\]